# SupplementSupplementary Information References A Fiber-pigtailed Quantum Dot Device Generating Indistinguishable Photons at GHz Clock-rates

### v1

@@ Abstract
Solid-state quantum light sources based on semiconductor quantum dots (QDs) are increasingly employed in photonic quantum information applications. Especially when moving towards real-world scenarios outside shielded lab environments, the efficient and robust coupling of nanophotonic devices to single-mode optical fibers offers substantial advantage by enabling "plug-and-play" operation. In this work we present a fiber-pigtailed cavity-enhanced source of flying qubits emitting single indistinguishable photons at clock-rates exceeding GHz. This is achieved by employing a fully deterministic technique for fiber-pigtailing optimized QD-devices based on hybrid circular Bragg grating (hCBG) micro-cavities. The fabricated fiber-pigtailed hCBGs feature radiative emission lifetimes of ps, corresponding to a Purcell factor of 9, a suppression of multi-photon emission events with %, a photon-indistinguishability % and a measured single-photon coupling efficiency of 53% in a high numerical aperture single-mode fiber, corresponding to 1.2 Megaclicks per second at the single-photon detectors. Our results show that fiber-pigtailed quantum light sources based on hCBG cavities are a prime candidate for applications of quantum information science.
@@ 1 Introduction
In this work, we report a fully deterministic fiber-pigtailed quantum light source based on QD-hCBG cavities operated at GHz clock-rates. The device exhibits -times as low as 76 ps, corresponding to a Purcell factor close to 9. We observe a strong suppression of multi-photon emission events associated with at 80 MHz excitation rate and pulsed two-photon indistinguishabilities of up to 82(4)% (79(4)%) for 2 ns (12.5 ns) temporal separation of consecutively emitted photons under pulsed p-shell-resonant excitation. The device features a single-photon in-fiber efficiency per excitation pulse of up to 53.7(2)%, corresponding to 1.2 Million clicks per second at the single-photon detectors. Furthermore we demonstrate that the Purcell-enhanced allows for driving the fiber-pigtailed device at 1.28 GHz clock-rate, providing single indistinguishable photons (%) at application-relevant GHz clock-rates.
@@ 2 Device Fabrication
Deterministic fiber-pigtailing of individual hCBG devices is achieved by employing the process outlined in the following. In a first step, the fiber glued into a ceramic ferrule is aligned to the target photonic structure with an interferometric positioning procedure: The sample surface is illuminated with a supercontinuum source through this fiber placed at a few hundred nanometer distance from the sample. The interference of light reflected from the sample and partially reflected from the fiber’s end face is recorded on a spectrometer. The spectral interference signal is used as feedback to adjust the tilt and height of the fiber with respect to the sample surface plane, and to locate the targeted hCBG structure while scanning the fiber across the sample surface. The precision of fiber-alignment to the targeted hCBG is expected to be below 200 nm. Next, the fiber facet is placed in physical contact with the sample to ensure stability. To achieve a specific distance between the fiber facet and the hCBG, the fiber-facet is etched prior the coupling procedure within an area of m around the fiber-core using a focused Xenon ion beam. In this way, the distance between the target QD-hCBG cavity and the fiber-core can be controlled with a precision of 50 nm. In the final fiber-pigtailing step, UV-sensitive adhesive is applied around the fiber-ferrule encapsulating the UHNA3 SMF and cured using a focused UV light source. The FC-device is then transferred to an optical setup for characterization. For details we refer to Ref. [37], where this method was applied to QD mesa structures.
@@ 3 FEM Simulations
Fig. 1(b) shows the simulated optical properties in terms of Purcell factor , free-space performance of the QD-hCBG cavity into a lens without fiber and simulated dipole power coupled to the UHNA3 fiber’s two degenerate ground modes. For =400 nm, simulated -values exceed 30, while direct in-fiber efficiencies of % are reached. The in-fiber efficiency here corresponds to the coupled power in the two degenerate SMF modes, compared to total emitted power by the embedded dipole. If Purcell enhancement and lens-efficiency into a lens with NA 0.8 instead of the SMF is considered, the simulated QD-hCBGs performance in absence of the fiber is similar in terms of , with slightly higher efficiencies of . The lens-efficiency corresponds to the fraction of total dipole power collected in the far-field for the corresponding NA, accordingly.
The fiber-to-CBG distance has a significant influence on the achievable Purcell enhancement, as can be seen from Fig. 1(c) for a wavelength of nm. -values between 50 and 25 are observed in the simulations for 2001000 nm. This large influence stems from the close proximity of the fiber to the cavity, effectively becoming part of the cavity itself [31, 32] and affecting the -factor. Note that the drop in simulated at h200 nm originates from a shift in the cavity mode wavelength, since the -dependent simulation is carried out for a constant wavelength. The cavity mode starts to red-shift to longer wavelengths when the fiber comes in close proximitty to the hCBG structure, due to increased effective refractive index. The in-fiber efficiency is only marginally affected and stays between 80-85% over the investigated -range, indicating the narrow collimation of the hCBG-mode’s emission at these distances.
An additional parameter with a more significant effect on the in-fiber efficiency is the lateral displacement of the UHNA3’s fiber-core with respect to the center of the QD-hCBG cavity. The deviation from the maximum efficiency with lateral displacements of the fiber are shown in Fig. 1(d), with the indicated alignment accuracy of nm of the pigtailing process still yielding of the maximum fiber-efficiency.
@@ 4 Spectral Properties
In Figure 2, spectroscopic data of the selected QD-hCBG cavity before and after the fiber-pigtailing process is presented (sample temperature: =4.8 K). Fig. 2(a) depicts spectra of the fiber-pigtailed QD-hCBG emission after four consecutive cooldowns (FCCD) and the reference spectrum before fiber-pigtailing, each under off-resonant excitation at nm and excitation power 10-100 nW. The emission spectrum prior to the pigtailing (grey line) shows dominant positive trion states (multiple lines associated with excitonic X2+ and other higher hole states, the dominant X+ and two lines corresponding to biexcitonic XX+ transitions), commonly observed for the investigated InAs QDs and dependent on their background doping [39].
@@
The successful fiber-pigtailing is additionally confirmed by the emission spectra under quasi-resonant excitation in Fig. 2(b), revealing near-identical spectra throughout cooldown runs #2-4. Moreover, we observe wavelength shifts for the X+ state of +0.25 nm for cooldown #1 and -1.55 nm for the cooldowns #2-4 relative to the reference spectrum before fiber-coupling. The wavelength of the excitation laser at 3 W was set to the p-shell resonance for the X+ state, as identified from photoluminescence excitation (PLE) scans before and after pigtailing (see section). The PLE scans additionally confirm the observed wavelength shifts of the X+ transition by showing a corresponding shift in the of the p-shell, indicating that it is in fact a spectral shift resulting from a change of the QD’s spectral fingerprint rather than a measurement artifact (e.g. due to changes in the alignment of the detection path).
@@
Due to the fact that both the QD emission and the hCBG modes exhibited a red-shift during cooldown #1 compared to before the pigtailing, while the cooldowns afterwards showed a blue-shift, a temperature-induced shift is unlikely, not least as the proper thermalization of the fiber-pigtail was confirmed by comparing the temperature dependence of the emission intensity during the cooldown runs with theoretical predictions (see. S.I. section S3). We thus identified strain as the origin of the observed shifts in emission wavelengths, and also for the changes in dominant emission lines. A reasonable explanation could be that the induced strain affects the availability of excess holes required for a dominant X+ emission under off-resonant excitation, enabling excess electrons from nearby donors instead causing X- to be the dominant line. Under p-shell excitation, the X+ transition still remains dominant. The strain is likely built up due to the different thermal expansion coefficient of the hCBG substrate (i.e. GaAs, SiO2 and Au, see Fig. 1(a)) and fiber-pigtail (i.e. silica and UV-adhesive). The strain could also explain the shifts in mode wavelengths between the cooldowns, by changing the (effective) refractive index of the combined fiber-hCBG microcavity.
@@ 5 Efficiency and Purcell Enhancement
The measured -times of the X+ for the consecutive cooldowns of the pigtailed device are displayed in Fig. 3(d). The observed spectral mismatch of X+ wavelength and closest cavity mode and the measured cavity mode -factor as are additionally listed. The pigtailed device shows a considerably larger (X+) during cooldown #1, while interestingly also exhibiting a clearly increased compared to before the pigtailing. However, while the spectral mismatch between X+ and mode was -0.44(0.50) nm before, the pigtailed device showed a mismatch of +1.53(0.50) nm during the first cooldown. We attribute the longer -time to the larger spectral mismatch after the first cooldown. With the strain conditions settled after the first temperature cycle, the following cooldowns yield very similar times, alongside X+-mode mismatches between -0.62(0.50) nm and 0.15(0.50) nm. While the observed spectral mismatch is close to the value before the pigtailing, the -factors are clearly reduced for the cooldowns #2 to #4. We attribute the slightly higher times after the fiber-pigtailing to the decreased Purcell enhancement by the reduced caused by the distance dependent effect of the fiber on the here as predicted in Fig. 1(d). This indicates that the fiber-to-hCBG distance for the pitailed device is probably higher than the of 350 nm, which causes the observed slight decrease in compared to before. Note that the device does not reach the simulated of 30 and more already before the coupling, likely due to a non-ideal spatial integration of the QD.
@@
After confirming the single-photon nature of the fiber-pigtailed QD device’s emission, its efficiency can be estimated by analyzing the photon flux detected at the single photon detector in the experimental setup and taking into account transmission losses from the fiberpigtail to the detection system. The observed countrates before and after the pigtailing (in this case for cooldown #2) are listed le 1. A kcps (kilo clicks per second) is obtained for the X+ under p-shell excitation (W) on the SNSPDs, while the X- reaches Mcps at off-resonant excitation with nm (nW).
@@
By dividing the measured countrates from Table 1 by the excitation repetition rate of MHz, the end-to-end efficiency at the detectors is , yielding % and %. To compare these efficiencies to the simulated fiber-coupling efficiency in Fig. 1(a), we take the setup efficiency into account. Details about the efficiency estimation can be found in S.I., section S6. Considering the UHNA3-to-780HP splice , the fiber-feedthrough at the cryostat and transmission losses by the fiber-beamsplitter, several fiber-connectors, the spectrometer and the connection to the SNSPDs , we obtain a fiber-coupling efficiency into the pigtailed UHNA3 fiber of % and % for and , respectively. The observed fiber-coupling efficiency for X- is already surprisingly close to the simulated efficiency of %, c possible non-ideal lateral alignment between fiber and hCBG device,. Furthermore, the off-resonant excitation of the X- reduces the measured efficiency, since e.g. the neutral states are also excited (see cooldown #2 in Fig. 2(a)).Note here, that the high efficiency observed for the X- state besides it’s far detuning from the cavity modes highlights the potential of the broadband capabilities of hCBG devices also for applications bbenefitingfrom long radiative lifetimes. The large spectral cavity-emitter detuning substantially increases (X-), while the piber-coupling efficiency remains on a high level.
@@
The fiber-coupling efficiency of the X+ is significantly lower than expected, however this can be explained considering the fact, that the X+ was already limited in brightness prior to the pigtailing: Table 1 shows the observed peak-intensities at the CCD for the X+ under p-shell near saturation before the pigtailing with kcps, and with pigtailed fiber as kcps. Free-space and fiber-setup efficiencies to the CCD were comparable for the measurements, so that the ratio of peak-CCD counts for the X+ before and after the pigtailing can also act as a rough estimate for the fiber-coupling efficiency. yields , which is reasonably close to bearing the aforementioned experimental limitation in mind and thus further indicating a high degree of pigtailing precision. We observed, that the saturated X+ brightness for the incorporated QDs in the hCBGs before the pigtailing varied by up to a factor of 4 between QDs, which we attribute to changing dopant and defect layer environments, that provide the charge carries for the X+. The fiber-pigtailed device here was primarily chosen for the high Purcell enhancement of the X+, rather than the X+ brightness, and we plan to optimize both parameters in the future.
@@
The performance of the fiber-pigtailed QD device presented above compares favorably with an earlier report on CBG-based plug-and-play sources by Jeon et al. [34]. We achieve significantly higher and lower for the quasi-resonantly excited X+. The improvement in multiphoton-suppression we thereby attribute to the quasi-resonant excitation scheme implemented in our work. Moreover, our fully deterministic technology for QD-device integration and fiber-pigtailing enables us to increase the in-fiber efficiency of the X+ and especially X- state by more than a factor six. This advancement is further facilitated by the optimized hybrid back-reflector design in combination with the UHNA3-fiber and precise fiber-to-hCBG alignment accuracy.

### v2

# A Fiber-pigtailed Quantum Dot Device Generating Indistinguishable Photons at GHz Clock-rates

@@ Abstract
Solid-state quantum light sources based on semiconductor quantum dots (QDs) are increasingly employed in photonic quantum information applications. Especially when moving towards real-world scenarios outside shielded lab environments, the efficient and robust coupling of nanophotonic devices to single-mode optical fibers offers substantial advantage by enabling "plug-and-play" operation. In this work we present a fiber-pigtailed cavity-enhanced source of flying qubits emitting single indistinguishable photons at clock-rates exceeding GHz. This is achieved by employing a fully deterministic technique for fiber-pigtailing optimized QD-devices based on hybrid circular Bragg grating (hCBG) micro-cavities. The fabricated fiber-pigtailed hCBGs feature radiative emission lifetimes of ps, corresponding to a Purcell factor of 9, a suppression of multi-photon emission events with %, a photon-indistinguishability % and a measured single-photon coupling efficiency of 53% in a high numerical aperture single-mode fiber, corresponding to 1.2 Megaclicks per second at the single-photon detectors under 80 MHz excitation clock-rates. Furthermore, we show that high multi-photon suppression and indistinguishability prevail for excitation clock-rates exceeding GHz. Our results show that Purcell-enhanced fiber-pigtailed quantum light sources based on hCBG cavities are a prime candidate for applications of quantum information science.
@@ 1 Introduction
In this work, we report a fully deterministic fiber-pigtailed quantum light source based on QD-hCBG cavities exhibiting -times as low as 76 ps, corresponding to a Purcell factor close to 9, which allows for operation at GHz clock-rates. We observe a strong suppression of multi-photon emission events associated with at 80 MHz excitation rate and pulsed two-photon indistinguishabilities of up to 82(4)% (79(4)%) for 2 ns (12.5 ns) temporal separation of consecutively emitted photons under pulsed p-shell-resonant excitation. The device features a single-photon fiber-coupling efficiency per excitation pulse of up to 53.7(2)%, corresponding to 1.2 Million clicks per second at the single-photon detectors at 80 MHz excitation frequency. Furthermore we demonstrate that the Purcell-enhanced allows for driving the fiber-pigtailed device at 1.28 GHz excitation clock-rate, providing single indistinguishable photons (%) at application-relevant GHz clock-rates.
@@ 2 Device Fabrication
Deterministic fiber-pigtailing of individual hCBG devices is achieved by employing the process outlined in the following. In a first step, the fiber glued into a ceramic ferrule is aligned to the target photonic structure with an interferometric positioning procedure: The sample surface is illuminated with a supercontinuum source through this fiber placed at a few hundred nanometer distance from the sample. The interference of light reflected from the sample and partially reflected from the fiber’s end face is recorded on a spectrometer. The spectral interference signal is used as feedback to adjust the tilt and height of the fiber with respect to the sample surface plane, and to locate the targeted hCBG structure while scanning the fiber across the sample surface. The precision of fiber-alignment to the targeted hCBG is expected to be below 200 nm. Next, the fiber facet is placed in physical contact with the sample to ensure stability. To achieve a specific distance between the fiber facet and the hCBG, the fiber-facet is etched prior the coupling procedure within an area of m around the fiber-core using a focused Xenon ion beam. In this way, the distance between the target QD-hCBG cavity and the fiber-core can be controlled with a precision of 50 nm. In the final fiber-pigtailing step, UV-sensitive adhesive is applied around the fiber-ferrule encapsulating the UHNA3 SMF and cured using a focused UV light source. The FC-device is then transferred to an optical setup for characterization. For details on the fiber splicing, alignment, and etching as well as investigations on the coupling stability, we refer to Ref. [67], applying the methods used in this work to QD mesa structures.
@@ 3 FEM Simulations
Fig. 1(b) shows the simulated optical properties in terms of Purcell factor , free-space performance of the QD-hCBG cavity into a lens without fiber and simulated dipole power coupled to the UHNA3 fiber’s two degenerate ground modes. For a target value of = 350 nm, simulated -values exceed 30, while direct single photon fiber-coupling efficiencies of are reached. The experimentally obtained precision of the distance between the QD-hCBG cavity and the fiber core is 50 nm [37], which is marked in Fig. 1(c) with dashed lines. This error margin corresponds to the change of the -values in the range of 10-20% and has a negligible effect on . The simulated here corresponds to the coupled power in the two degenerate SMF modes, compared to total emitted power by the embedded dipole, which is equal to the probability of collecting a single photon per excitation pulse, assuming unity QD quantum efficiency and preparation fidelity of the excitation. If Purcell enhancement and efficiency into a lens with NA 0.8 instead of the SMF is considered, the simulated QD-hCBGs performance in absence of the fiber is similar in terms of , with slightly higher efficiencies of . The lens-efficiency corresponds to the fraction of total dipole power collected in the far-field for the corresponding NA, accordingly.
The fiber-to-CBG distance has a significant influence on the achievable Purcell enhancement, as can be seen from Fig. 1(c) for a wavelength of nm. -values between 50 and 25 are observed in the simulations for nm. This large influence stems from additional vertical confinement to the hCBG mode in vertical direction by the present fiber. As shown in S.I., section S2, placing the fiber-facet at regions of high (low) hCBG mode intensity in vertical direction, causes a decrease (increase) in the mode’s -factor by up to 45%. Since the mode volume stays nearly constant, the -dependency of leads to an -dependent . The fiber can be considered as effectively becoming part of the hCBG cavity. Noteworthy, the -dependent should allow to access the fiber-to-hCBG distance via mode linewidth measurements in the experiment. Note that the drop in simulated at h200 nm originates from a shift in the cavity mode wavelength, since the -dependent simulation is carried out for a constant wavelength. The cavity mode starts to red-shift to longer wavelengths when the fiber comes in close proximity to the hCBG structure, due to increased effective refractive index. is only marginally affected and stays between 80-85% over the investigated -range, indicating the narrow collimation of the hCBG-mode’s emission at these distances.
An additional parameter with a more significant effect on the is the lateral displacement of the UHNA3’s fiber-core with respect to the center of the QD-hCBG cavity. The deviation from the maximum efficiency with lateral displacements of the fiber are shown in Fig. 1(d), with the indicated alignment accuracy of nm of the pigtailing process still yielding of the maximum fiber-efficiency.
@@ 4 Spectral Properties
In Figure 2, spectroscopic data of the selected QD-hCBG cavity before and after the fiber-pigtailing process is presented (sample temperature: =4.8 K). Fig. 2(a) depicts spectra of the fiber-pigtailed QD-hCBG emission after four consecutive cooldowns (FCCD) and the reference spectrum before fiber-pigtailing, each under off-resonant excitation at nm and excitation power 10-100 nW. The emission spectrum prior to the pigtailing (grey line) shows dominant positive trion states (multiple lines associated with excitonic X2+ and other higher hole states, the dominant X+ and two lines corresponding to biexcitonic XX+ transitions), commonly observed for the investigated InAs QDs and dependent on their background doping [49]. The assignment of the emission lines to specific QD states was confirmed by conducting excitation power- and polarization-resolved measurements after the first cooldown of the fiber-pigtailed device (see S.I., section S5. for details).
@@
The successful fiber-pigtailing is additionally confirmed by the emission spectra under quasi-resonant excitation in Fig. 2(b), revealing near-identical spectra throughout cooldown runs #2-4. Moreover, we observe wavelength shifts for the X+ state of +0.25 nm for cooldown #1 and -1.55 nm for the cooldowns #2-4 relative to the reference spectrum before fiber-coupling. The wavelength of the excitation laser at 3 W was set to the p-shell resonance for the X+ state, as identified from photoluminescence excitation (PLE) scans before and after pigtailing (see S.I. section S5). The PLE scans additionally confirm the observed wavelength shifts of the X+ transition by showing a corresponding shift in the of the p-shell, indicating that it is in fact a spectral shift resulting from a change of the QD’s spectral fingerprint rather than a measurement artifact (e.g. due to changes in the alignment of the detection path).
@@
Due to the fact that both the QD emission and the hCBG modes exhibited a red-shift during cooldown #1 compared to before the pigtailing, while the cooldowns afterwards showed a blue-shift, a temperature-induced shift is unlikely, not least as the proper thermalization of the fiber-pigtail was confirmed by comparing the temperature dependence of the emission intensity during the cooldown runs with theoretical predictions (see. S.I. section S3). We thus identify strain as the origin of the observed shifts in emission wavelengths and also the changes in dominant emission lines. A reasonable explanation could be that the induced strain affects the availability of excess holes required for a dominant X+ emission under off-resonant excitation, enabling excess electrons from nearby donors instead causing X- to be the dominant line. Under p-shell excitation, the X+ transition still remains dominant. The strain is likely built up due to the different thermal expansion coefficient of the hCBG substrate (i.e. GaAs, SiO2 and Au, see Fig. 1(a)) and fiber-pigtail (i.e. silica and UV-adhesive). The strain could also explain the shifts in mode wavelengths between the cooldowns, by changing the (effective) refractive index of the combined fiber-hCBG microcavity. For a rough estimation based on FEM simulations, we find that a refractive index change 0.2% is already sufficient to create the observed mode shifts. Such a is well within reach for already small amounts of strain [55].
@@ 5 Efficiency and Purcell Enhancement
The measured -times of the X+ for the consecutive cooldowns of the pigtailed device are displayed in Fig. 3(d). The observed spectral mismatch of X+ wavelength and closest cavity mode and the measured cavity mode -factor as are additionally listed. The pigtailed device shows a considerably larger (X+) during cooldown #1, while interestingly also exhibiting a clearly increased compared to before the pigtailing. However, while the spectral mismatch between X+ and mode was -0.44(0.50) nm before, the pigtailed device showed a mismatch of +1.53(0.50) nm during the first cooldown. We attribute the longer -time to the larger spectral mismatch after the first cooldown. With the strain conditions settled after the first temperature cycle, the following cooldowns yield very similar times, alongside X+-mode mismatches between -0.62(0.50) nm and 0.15(0.50) nm. While the observed spectral mismatch is close to the value before the pigtailing, the -factors are clearly reduced for the cooldowns #2 to #4. We attribute the slightly higher times after the fiber-pigtailing to the decreased Purcell enhancement by the reduced caused by the distance-dependent effect of the fiber on the here as predicted in Fig. 1(dc).
The decrease in experimental between before the pigtailing and for cooldown #1 is approximately , which is also close to the relation between the simulated before the pigtailing and at nm of 430/320, considering that the overall was reduced slightly through imperfect fabrication. This indicates that before cooldown #1, the distance of fiber and hCBG cavity was potentially close to the target value. After the warmup and following cooldowns #2 to #4, the experimental drops to 0.71 the -factor before the pigtailing, which means that the fiber-to-hCBG distance for the pitailed device afterwards might be up to 200 nm larger than the of 350 nm. We note that the expected strain-related refractive index change causing the spectral shift of the modes results in a negligible change in as confirmed from the simulations, hinting further towards a fiber-distance related effect. The resulting -times before and after the pigtailing during cooldowns #2 to #4 exhibit a similar relation of 0.72, implying that the change in could stem from the distance-dependent reduced by the lower . Future work simulating the strain conditions [45] could answer the question of how much is additionally affected by the strain itself. Note that the device does not reach the simulated of 30 and more already before the coupling, likely due to a non-ideal spatial integration of the QD.
@@
After confirming the single-photon nature of the fiber-pigtailed QD device’s emission, its efficiency can be estimated by analyzing the photon flux detected at the single photon detector in the experimental setup and taking into account transmission losses from the fiberpigtail to the detection system. The observed countrates before and after the pigtailing (in this case for cooldown #2) are listed in Table 1. A countrate of kcps (kilo clicks per second) is obtained for the X+ under p-shell excitation (W) on the SNSPDs, while the X- reaches Mcps at off-resonant excitation with nm (nW).
@@
By dividing the measured countrates from Table 1 by the excitation repetition rate of MHz, the end-to-end efficiency at the detectors is , yielding % and %. To compare these efficiencies to the simulated in Fig. 1(a), we take the setup efficiency into account. Details about the efficiency estimation can be found in S.I., section S8. Considering the UHNA3-to-780HP splice , the fiber-feedthrough at the cryostat and transmission losses by the fiber-beamsplitter, several fiber-connectors, the spectrometer and the connection to the SNSPDs , we obtain an experimental into the pigtailed UHNA3 fiber of% and % for and , respectively.
The observed for X- is already surprisingly close to the simulated efficiency of %, a possible non-ideal lateral alignment between fiber and hCBG device. Furthermore, the off-resonant excitation of the X- reduces the measured efficiency, since e.g. the neutral states are also excited (see cooldown #2 in Fig. 2(a)). Note here, that the high efficiency observed for the X- state besides it’s far detuning from the cavity modes highlights the potential of the broadband capabilities of hCBG devices also for applications benefiting from long radiative lifetimes. The large spectral cavity-emitter detuning substantially increases (X-), while remains on a high level.
@@
The observed textcolorred of the X+ is significantly lower than expected, however this can be explained considering the fact, that the X+ was already limited in brightness prior to the pigtailing: Table 1 shows the observed peak-intensities at the CCD for the X+ under p-shell near saturation before the pigtailing with kcps, and with pigtailed fiber as kcps. Free-space and fiber-setup efficiencies to the CCD were comparable for the measurements, so that the ratio of peak-CCD counts for the X+ before and after the pigtailing can also act as a rough estimate for . yields , which is reasonably close to the simulated bearing the aforementioned experimental limitation in mind and thus further indicating a high degree of pigtailing precision. We observed, that the saturated X+ brightness for the incorporated QDs in the hCBGs before the pigtailing varied by up to a factor of 4 between QDs, which we attribute to changing dopant and defect layer environments, that provide the charge carries for the X+. The fiber-pigtailed device here was primarily chosen for the high Purcell enhancement of the X+, rather than the X+ brightness, and we plan to optimize both parameters in the future.
@@
The performance of the fiber-pigtailed QD device presented above compares favorably with an earlier report on CBG-based plug-and-play sources by Jeon et al. [23]. We achieve significantly higher and lower for the quasi-resonantly excited X+. The improvement in multiphoton-suppression we thereby attribute to the quasi-resonant excitation scheme implemented in our work. Moreover, our fully deterministic technology for QD-device integration and fiber-pigtailing enables us to increase of the X+ and especially X- state by more than a factor six. This advancement is further facilitated by the optimized hybrid back-reflector design in combination with the UHNA3-fiber and precise fiber-to-hCBG alignment accuracy.